# Bidirectional Uncertainty-Aware Region Learning for Semi-Supervised Medical Image Segmentation

### v1

@@ 4.5.2 Regional Learning Strategies
4.5.3 Effect of Uncertainty Thresholds $\mu$ on Model Performance
In this section, we explore the effect of the uncertainty threshold $\mu$ on model performance. Specifically, choosing the 95th percentile to identify uncertainty regions in labeled data aims to comprehensively cover areas where errors may occur and rely on accurate labels to correct them. Conversely, selecting the 99th percentile in unlabeled data aims to more accurately pinpoint uncertainty regions. As shown in Fig. 5, When the $\mu$ is set at the 95th and 97th percentiles, some uncertainty regions partially appear in the background. Generally speaking, background features are relatively uniform and consistent, and the model can receive many background samples during training, resulting in a lower probability of prediction errors in background regions. Therefore, the choice of the 99th percentile aims to avoid misclassifying correct regions as uncertainty regions while maximizing the precision of uncertainty region localization. To validate this hypothesis, we conducted the relevant experiments, as shown in Table 5: Compared to the baseline, the combinations of $\mu$ set at 95%+95% and 99%+99% for labeled and unlabeled data both exhibit reliable performance, indirectly validating the effectiveness of the proposed module. However, their performance is still lacking compared to the 95%+99% combination. This suggests that the mixed approach of using a broader uncertainty range (95%) in labeled data and a more precise uncertainty selection (99%) in unlabeled data achieves the optimal balance. This combination effectively enhances the model’s robustness, providing ample correction opportunities for labeled data while avoiding misjudgment of uncertain regions in unlabeled data.
Additionally, in the initial stage of training with unlabeled data, we visualize the pseudo-labels generated by the teacher model. As shown in Fig. 6, in BCP, the pseudo-labels generated by the teacher model contain many errors. However, in Ours(BCP), the pseudo-label errors generated by the teacher model trained with the 99th percentile threshold are slightly reduced. In contrast, those generated with the 95th percentile threshold are significantly fewer. This demonstrates the feasibility of the settings in our approach.
@@ 4.5.4 Remove Unreliable Predictions on Model Learning
In addition, we investigate the impact of directly removing unreliable predictions during model training in Our(BCP) with 10% labeled data. As shown in Tabel 6. Experimental results demonstrate that even erroneous prediction regions can contribute to model training. Therefore, we can reduce the learning weights of these predicted regions instead of completely removing them.
@@ 5 Conclusion
In this paper, we propose a simple semi-supervised training method focusing on uncertainty region learning for medical image segmentation. Our method employs a bidirectional uncertainty-aware region Learning strategy. First, in the labeled data phase, we focus on processing the uncertainty regions to improve the model’s segmentation accuracy of the uncertainty regions through precise labeling information. Second, in the unlabeled data phase, since pseudo-labeling errors cannot be avoided, we adjust the learning focus of the model by calculating the uncertainty region so that it focuses on learning the deterministic region and reduces the learning weight on the uncertainty region, which can mitigate the learning interferences introduced by erroneous pseudo-labeling. We demonstrate better performance than current state-of-the-art methods on popular datasets (LA and ACDC). Given its simplicity and superior performance, we believe that this method deserves to be further explored in future studies to understand its applicability across different datasets and domains.

### v2

@@ 4.5.5 Remove Unreliable Predictions on Model Learning
We further investigate the effect of directly removing unreliable predictions on model performance when training the model on the ACDC dataset with 10% labeled data. As shown in Table 8, the experimental results show that simply removing these unreliable predictions instead leads to a decrease in model performance. This suggests that even incorrectly predicted regions can still provide valuable training information for the model to some extent. Completely discarding them not only reduces the amount of available training data but may also hinder the model’s ability to adapt to complex or ambiguous regions. In contrast, our method retains these regions and adjusts their influence through adaptive weighting, allowing the model to benefit from difficult samples while minimizing the risk of being misled by noisy supervision. This strategy strikes a better balance between leveraging information and ensuring training stability.
Dataset ACDC LA
Method BCP Ours BCP Ours
FLOPs 3.49G 3.53G 6.00G 6.16G
Runtime 3.15 h 3.52 h 9.67 h 10.25 h
Table 9: Comparison of training cost between BCP and our method.
4.5.6 Runtime and Efficiency
We compare the FLOPs and runtime between BCP and our method. As shown in Table 9, our method introduces only a slight increase time, primarily attributed to the lightweight entropy-based computation and region-weighting strategy. These additional operations incur minimal computational overhead and do not substantially affect overall training efficiency, demonstrating the practical feasibility of our approach for real-world clinical deployment.
5 Limitations and Future Directions
Although our proposed bi-directional uncertainty-aware region learning strategy has demonstrated significant performance advantages in several medical image segmentation tasks, especially in improving the model’s ability to perceive difficult regions and attenuating pseudo-labelling misrepresentation, there are still some potential limitations of the method that deserve further exploration. Firstly, the division of uncertainty regions relies on fixed percentile thresholds (e.g., 95% and 99%, which may need to be readjusted to suit their characteristics when facing different datasets or more complex tasks. Second, the learning weight adjustment strategy we impose on uncertain regions during training is still an empirical setting. That is, enhancing the learning intensity of high uncertainty regions in labelled data and weakening the learning intensity of uncertainty regions in unlabelled data, this ‘raise-lower’ mechanism is intuitive and effective, but lacks the ability of more fine-grained dynamic adjustment. Therefore, future research can further explore more adaptive mechanisms to improve model stability and generalisation across datasets.
@@ 6 Conclusion
In this paper, we propose a semi-supervised training method based on a two-way uncertainty-aware region learning strategy, which employs different uncertainty region learning strategies for labeled and unlabeled data, to enhance the model’s ability to learn the samples from the difficult regions, and at the same time, reduce the negative impact of these samples on the training of the model. Specifically, for the labeled data, we focus on processing the uncertainty regions to improve the model’s segmentation accuracy of the uncertainty regions through precise labeling information. Second, for the unlabeled data, since pseudo-labeling errors cannot be avoided, we adjust the learning focus of the model so that it focuses on learning the deterministic region, which can mitigate the learning interferences introduced by erroneous pseudo-labeling. With the collaboration of these two strategies, the model can better leverage the characteristics of both labeled and unlabeled data, achieving more stable and precise segmentation performance.
Declarations
Funding This study is funded by the National Natural Science Foundation of China (62076005, U20A20398), the Natural Science Foundation of Anhui Province (2008085MF191), the University Synergy Innovation Program of Anhui Province, China (GXXT-2021-002), and the Provincial Quality Project of Education in the New Era in 2023 (Postgraduate Education 2023lhpysfjd009).
Data Availability The data that support the findings of this study are available from the corresponding author upon reasonable request.
Confict of Interest The authors declare that they have no conflict of interest.